For the system to do 2000 J work, 2500 J of heat must be added to it. Calculate 

the possible changes in the internal energy of the system.



1
Expert's answer
2021-06-13T11:29:28-0400
"\Delta U=Q-W\\\Delta U=2500-2000=500\ J"